Two similar charges of 250 static are situated on small spheres 15 cm apart. What is the electric field intensity midway between the spheres? What is the potential at this point?


1
Expert's answer
2021-05-13T07:19:30-0400

work done = force "\times distance = Nm" or J


Force = k"\frac{q_1q_2}{r^2}= 9\times 10^9 \times \frac{250\times 10^{-6}\times 250 \times 10^{-6}}{0.15^2}"


force = 2500 N

Electric potential energy PE= force"\times distance"

PE= "2500 \times 0.015= 3750 J"


Electric potential = "k\times \frac{Q}{r}" = "\frac{3750}{0.15}\times 9.0 \times 10^9 = 2.14 \times 10^{14} J/C"